OpenStudy (anonymous):

Computer Capacity is often measured in bits and bytes. A bit is the smallest unit, which is 1 or 0, in the computers memory. A byte is 2^3 bits. a megabyte is 2^20 bytes. How many bytes are in a megabyte?

OpenStudy (anonymous):

1048576 bytes in a megabyte.

OpenStudy (anonymous):

did you just multipy the powers or did you add them?

OpenStudy (anonymous):

Well, you asked how many bytes in a megabyte, so I just gave you the answer you had in the problem? If you want BITS in a megabyte, its: 8388608 which is (2^3) * (2^20) bits in a byte.

OpenStudy (anonymous):

oh so all i had to do was use 2^20?

OpenStudy (anonymous):

@satellite73 can you help me? idk how he got the answer or if its even right

OpenStudy (anonymous):

Givens: \[1 = \frac{2^3 bits}{byte}\]\[1 = \frac{10^{20} MB}{byte}\] Therefore, if your question is how many bytes in a megabyte, the answer is\[10^{20}\]. However, if you're looking for the number of bits in a megabyte:\[1 = \frac{2^3 bits}{byte} * \frac{10^{20} bytes}{megabyte} = \frac{8 * 10^{20} bits}{megabyte}\] Therefore, the answer would be \[8*10^{20} \frac{bits}{megabyte}\]

OpenStudy (anonymous):

oh ok. now i get it. i just didnt know how he just had that answer